Find lim x→0 (1-cos2x) / (x2 sinx2)



1
Expert's answer
2021-05-31T15:32:36-0400

Let us find the limit "\lim\limits_{ x\to 0} \frac{1-\cos^2x}{ x^2 \sin x^2}:"


"\lim\limits_{ x\to 0} \frac{1-\cos^2x}{ x^2 \sin x^2}=
\lim\limits_{ x\to 0} \frac{\sin^2x}{ x^2 \sin x^2}=
\lim\limits_{ x\to 0} \frac{\sin x}{ x}\cdot \lim\limits_{ x\to 0} \frac{\sin x}{ x}\cdot \lim\limits_{ x\to 0} \frac{1}{ \sin x^2}=1\cdot 1\cdot \infty=\infty."